OpenStudy (anonymous):

What is the perimeter of a rectangular driveway that measures 25 feet by 60 feet? Choose from one of these:170,210,250,1500

OpenStudy (anonymous):

Need help

OpenStudy (anonymous):

the perimeter is 2L+2W so 2(60)+2(25) = 120+50=170

OpenStudy (anonymous):

where L and W are length and width

OpenStudy (mathstudent55):

The perimeter of a figure is the sum of the lengths of all sides. In the case of a rectangle, you have 4 sides. Two sides are lengths and two sides are widths. Add them all up, and you'll get the answer.

OpenStudy (anonymous):

So the answer is 170

OpenStudy (anonymous):

Thank you